# Convolutional autoencoders for the reconstruction of three-dimensional interfacial multiphase flows

### v1

@@ 3.6 Generalization error
Finally, we investigate the generalization ability of each interface representation by evaluating the mean Dice coefficient for each combination of train/test dataset. Figure 11 shows that sharper interface representations transfer better to out-of-domain test datasets. Furthermore, sharp interface models trained only on synthetic data transfer surprisingly effectively to the HIT dataset, suggesting that synthetic data augmentation is an effective tool in this domain.
Figure 11: Comparison of generalization error for different interface representations when training and evaluating on different combinations of datasets.
@@ 4 Conclusions
We have evaluated the performance of a fully convolutional 3D autoencoder in reconstructing interfacial geometries using common implicit interface representations employed in multiphase simulations. Our comparisons across simulated and synthetic datasets of varying complexity revealed that the signed distance function consistently yields poorer reconstruction performance. In contrast, sharp and diffuse interface representations exhibit a trade-off: sharper interfaces improve the reconstruction of large-scale features, but excessively sharp representations tend to miss small-scale structures. On the other hand, overly diffuse interfaces lead to degraded overall accuracy.
Overall, we find that a moderately diffuse interface representation, with an interface width of 141-4 times the grid spacing, offers the best balance across datasets and metrics, preserving small features while maintaining high reconstruction fidelity.
Future work will focus on improving the autoencoder by incorporating conservation constraints on the network output, exploring variational autoencoders, and including additional state variables relevant to multiphase flows. Moreover, the insights gained from this comprehensive analysis lay the groundwork for developing models capable of predicting temporal dynamics in multiphase systems.
@@ Acknowledgements
Appendix A Best performing hyper-parameter sets

### v2

@@ 3.5 Generalization error
Finally, we investigate the generalization ability of each interface representation by evaluating the mean Dice coefficient for each combination of train/test dataset. For computational reasons, only 10% of each test was used, so the absolute values of the dice coefficient are not directly comparable to Figure 6. Figure 9 shows that sharper interface representations transfer better to out-of-domain test datasets. Furthermore, sharper interface models trained only on synthetic data transfer surprisingly effectively to the HIT dataset, suggesting that synthetic data augmentation is an effective tool in this domain. Moreover, the SDF is not only the least accurate, but also suffers from training instabilities for the most complex synthetic dataset ($\mu=1$). Specifically, the row of zeros in bottom right subfigure indicates a lack of training convergence.
Figure 9: Comparison of generalization error for different interface representations when training and evaluating on different combinations of datasets. Note that zero dice coefficient indicates that training did not converge.
@@ 4 Conclusions
We have evaluated the performance of a fully convolutional three-dimensional autoencoder for reconstructing interfacial geometries using common implicit interface representations employed in multiphase simulations. Comparisons across simulated and synthetic datasets of varying complexity show that the signed-distance function consistently yields inferior reconstruction performance. In contrast, sharp and diffuse interface representations exhibit a clear trade-off: sharper interfaces improve the reconstruction of large-scale features, but excessively sharp representations tend to miss small-scale structures, while overly diffuse interfaces degrade overall accuracy.
Across all datasets and metrics considered, a moderately diffuse interface representation with an interface width of one to four grid spacings provides the best balance, preserving small-scale features while maintaining high reconstruction fidelity. We further demonstrate that this trade-off can be explained by the spectral bias of convolutional neural networks. Spectral analysis of the reconstructed fields confirms that high-wavenumber content associated with sharp interfaces is systematically under-represented, leading to increased errors at small scales. Moderately diffuse interfaces mitigate this effect by regularizing high-frequency content, resulting in more accurate reconstructions across spatial scales.
Due to the high computational cost of training three-dimensional autoencoders, this study considered a single compression ratio and a single architecture type, and training uncertainty was only partially quantified. Future work will focus on incorporating conservation constraints into the network output, exploring variational autoencoders, and including additional state variables relevant to multiphase flows. The insights gained from this analysis lay the groundwork for developing reduced-order models capable of predicting temporal dynamics in complex multiphase systems.
@@ Acknowledgements
Appendix A Model architecture
See table 3.
Table 3: Detailed network architecture. An input volume of shape $1\times H\times W\times D$ is processed. ResBlocks contain SiLU activations, and all convolutions use padding to preserve dimensions unless strided. The final decoder layer is linear. $Z$ denotes the number of latent channels (default is 4).
Component Layer Output Shape
Encoder
Input - $1\times H\times W\times D$
Initial Conv Conv3d(1, 32) $32\times H\times W\times D$
Stage 1 2 $\times$ ResBlock(32) $32\times H\times W\times D$
Downsample Conv3d(32, 64) $64\times\frac{H}{2}\times\frac{W}{2}\times\frac{D}{2}$
Stage 2 2 $\times$ ResBlock(64) $64\times\frac{H}{2}\times\frac{W}{2}\times\frac{D}{2}$
Downsample Conv3d(64, 128) $128\times\frac{H}{4}\times\frac{W}{4}\times\frac{D}{4}$
Stage 3 2 $\times$ ResBlock(128) $128\times\frac{H}{4}\times\frac{W}{4}\times\frac{D}{4}$
Downsample Conv3d(128, 256) $256\times\frac{H}{8}\times\frac{W}{8}\times\frac{D}{8}$
Stage 4 2 $\times$ ResBlock(256) $256\times\frac{H}{8}\times\frac{W}{8}\times\frac{D}{8}$
Downsample Conv3d(256, 256) $256\times\frac{H}{16}\times\frac{W}{16}\times\frac{D}{16}$
Bottleneck 4 $\times$ ResBlock(256) $256\times\frac{H}{16}\times\frac{W}{16}\times\frac{D}{16}$
Conv3d(256, 256) $256\times\frac{H}{16}\times\frac{W}{16}\times\frac{D}{16}$
Final Encoder Conv Conv3d(256, Z) $Z\times\frac{H}{16}\times\frac{W}{16}\times\frac{D}{16}$
Decoder
Initial Decoder Conv Conv3d(Z, 256) $256\times\frac{H}{16}\times\frac{W}{16}\times\frac{D}{16}$
Stage 1 2 $\times$ ResBlock(256) $256\times\frac{H}{16}\times\frac{W}{16}\times\frac{D}{16}$
Upsample ConvTranspose3d(256, 256) $256\times\frac{H}{8}\times\frac{W}{8}\times\frac{D}{8}$
Stage 2 2 $\times$ ResBlock(256) $256\times\frac{H}{8}\times\frac{W}{8}\times\frac{D}{8}$
Upsample ConvTranspose3d(256, 128) $128\times\frac{H}{4}\times\frac{W}{4}\times\frac{D}{4}$
Stage 3 2 $\times$ ResBlock(128) $128\times\frac{H}{4}\times\frac{W}{4}\times\frac{D}{4}$
Upsample ConvTranspose3d(128, 64) $64\times\frac{H}{2}\times\frac{W}{2}\times\frac{D}{2}$
Stage 4 2 $\times$ ResBlock(64) $64\times\frac{H}{2}\times\frac{W}{2}\times\frac{D}{2}$
Upsample ConvTranspose3d(64, 32) $32\times H\times W\times D$
Final Blocks 4 $\times$ ResBlock(32) $32\times H\times W\times D$
Conv3d(32, 32) $32\times H\times W\times D$
Final Decoder Conv Conv3d(32, 1) $1\times H\times W\times D$
Appendix B Training loss curves
The training losses corresponding to the runs presented in Figure 6 are shown here in Figure 10. While the models have not reached complete convergence, the losses have largely plateaued, enabling meaningful comparison of their relative performance. Since the optimal learning rate was checked independently for each interface type through hyperparameter search, differences in convergence speed reflect the inherent learnability of each representation.
Figure 10: Training loss vs. epoch for all models compared in Figure 6
Appendix C Best performing hyper-parameter sets
@@
Appendix D Spectral analysis of reconstruction errors
To quantify how reconstruction accuracy varies across spatial scales, we compute radially averaged three-dimensional power spectral densities (PSDs) of the reconstructed interface fields and corresponding ground-truth samples. For each sample, the field is first mean-subtracted and transformed using a three-dimensional Fourier transform. The power spectrum is then obtained from the squared magnitude of the Fourier coefficients and radially averaged over spherical shells in wavenumber space. All spectra are computed consistently for the ground truth, the autoencoder reconstructions, and the reconstruction error (defined as the difference between reconstruction and ground truth), and are subsequently averaged over the test set. The analysis is performed for the most challenging synthetic dataset ($\mu=1$), where fine-scale interfacial structures are most pronounced. We focus on two representative interface formulations: a sharp interface representation and a diffuse interface representations based on a hyperbolic tangent profile with interface thicknesses $\epsilon=1/32$. Figure 12 shows the PSDs for the ground truth, reconstructions, and reconstruction errors for the two interface representations. Across both cases, the reconstructions closely match the ground-truth spectra at low wavenumbers. However, a systematic loss of energy is observed at high wavenumbers. This provides direct evidence that fine-scale features are preferentially lost in the reconstruction. Even in the ground truths (inputs to the autoencoder), the spectral content of the two interface representations are significantly different. The sharp interface representation exhibits a slow decay of energy content across wavenumbers. As a result, there is significant mismatch in the energy content of the reconstructed sharp interface fields (large errors) for moderate to high wavenumbers. On the other hand, the diffuse interface representations show strong decay of energy content across the wavenumbers. As such, the energy content of the error is much smaller than the sharp-interface representation for moderate to high wavenumbers (small scales). This analysis demonstrates the spectral bias of the autoencoder neural network, where the trained model struggles to reconstruct the small features and sharp interfaces that are dominant in the sharp interface representation.
(a) Sharp
(b) Tanh $1/32$
Figure 12: Radially averaged three-dimensional power spectral densities for the $\mu=1$ synthetic dataset. Results are shown for (a) sharp interface and (c) diffuse interface with $\epsilon=1/32$. In each case, the ground truth, reconstruction, and reconstruction error spectra are shown. The reconstructions systematically under-represent high-wavenumber content, with the severity of attenuation depending on the interface representation.
Appendix E Linear autoencoder baseline
To assess whether nonlinear representations are necessary for accurate reconstruction of interfacial multiphase flows, we consider a linear autoencoder baseline. A linear autoencoder consists of a single linear encoder and decoder without activation functions. When trained to optimality using a squared reconstruction loss, this model is mathematically equivalent to principal component analysis (PCA) [baldi1989neural]. Including this baseline allows us to evaluate whether the improved performance of convolutional autoencoders arises from their nonlinear feature extraction capabilities, rather than from architectural or optimization choices alone. The linear autoencoder is trained using the same datasets, train–validation splits, and compression ratio as the nonlinear convolutional autoencoder presented in the main text. In particular, the latent dimensionality is fixed to 256 for all experiments. The model is trained using the Adam optimizer to minimize the mean squared reconstruction error. While gradient-based optimization does not guarantee convergence to the global optimum of the linear model, the resulting performance provides a practical and relevant baseline for comparison under identical training conditions. Across all datasets and interface representations considered, the linear autoencoder exhibits substantially larger reconstruction errors than the nonlinear convolutional autoencoder. This performance gap is observed consistently for sharp, diffuse, and signed-distance interface representations. These results demonstrate that linear dimensionality reduction techniques such as PCA are insufficient for accurately representing complex three-dimensional interfacial geometries in multiphase flows, and confirm that nonlinear autoencoders provide a necessary and meaningful advantage for reduced-order representation in this setting.
Figure 13: Performance of each interface representation on the synthetic datasets, using a linear autoencoder model. See Figure 6 for corresponding results from the nonlinear convolutional autoencoder model.